# Enhancing Chemistry on Quantum Computers with Fermionic Linear Optical Simulation

### v1

@@ Abstract
We present and open source a simulator for circuits composed of passive fermionic linear optical elements and controlled-phase gates. Given such a circuit, our simulator can compute Born-rule probabilities for samples drawn from it. Our simulator supports both exact and approximate probability calculation, allowing users to trade accuracy for efficiency as needed. For approximate Born-rule probability calculation, our simulator’s runtime is exponential only in the magnitudes of the angles of the circuit’s controlled-phase gates. This makes our simulator useful for simulating certain systems that are beyond the reach of conventional state vector methods. We demonstrate our simulator’s utility by simulating the local unitary cluster Jastrow (LUCJ) ansatz and integrating it with sample-based quantum diagonalization (SQD) to improve the accuracy of molecular ground-state energy estimates. Applied to a 52-qubit N2 system, we observe accuracy improvements of up to 46% over the baseline SQD implementation with negligible computational overhead. As an efficient and flexible tool for simulating fermionic circuits, our simulator enables new opportunities for enhancing near-term quantum algorithms in chemistry and related domains.
@@ I Introduction
Among the most promising near-term applications of quantum computing is quantum computational chemistry [cao2019quantum, mcardle2020quantum], which has attracted considerable attention due to potentially transformative applications in drug development [santagati2024drug] and materials science [alexeev2024quantum], among other fields. Despite the excitement around this application, hardware demonstrations have been limited by the noise present in current quantum processors. Quantum error correction [gottesman1997stabilizer, fowler2012surface, shor1995scheme, steane1996error, calderbank1996good, kitaev2003fault] offers a long-term solution to faulty hardware, but a fully fault-tolerant quantum computer has yet to be built. In the near-term, quantum error mitigation techniques [temme2017error, li2017efficient, huggins2021virtual, liao2024machine, cincio2021machine] can improve results obtained from noisy quantum processors, in which many of these techniques involve classical simulation of quantum circuits [czarnik2021error, filippov2023scalabletensornetworkerrormitigation, eddins2024lightcone]. Because classical simulation of quantum circuits is intractable in general, these simulations must introduce approximations, or are exact only when applied to specific narrow classes of circuits. Simulation methods tailored for circuits relevant to quantum chemistry may enable new techniques for enhancing the performance of chemistry workloads on quantum computers.
@@
In this work, we introduce ExtraFerm [^1] https://github.com/zhassman/ExtraFerm, an open-source quantum circuit simulator that can be used to compute the Born-rule probabilities of samples (bitstrings) from circuits composed of passive fermionic linear optical elements and controlled-phase gates. These circuits are highly relevant to the field of quantum computational chemistry, in which passive fermionic linear optical elements are also known as particle number-conserving matchgates. ExtraFerm can perform both exact and approximate probability calculations for measurements in the output distributions of these circuits. For approximate calculation, the only exponential component of ExtraFerm’s runtime is a linear dependence in a quantity called the extent of the circuit. The extent is larger for controlled phase gates with larger phases and is multiplicative as more controlled phase gates are added (see Refs. [reardon2024improved, Dias2024classicalsimulation, reardonsmith2024fermioniclinearopticalextent]). For exact probability calculation, the runtime is exponential only in the number of controlled-phase gates.
@@
Notably, this means that for both approximate and exact calculation, ExtraFerm’s performance is polynomial in the number of qubits and matchgates. Unlike conventional state vector simulation, which tracks the amplitudes of all bitstrings, ExtraFerm computes Born-rule probabilities for a pre-specified subset of the output distribution. Indeed, for small systems, one can use ExtraFerm to simulate the entire support of a circuit, but we do not expect this to be the most compelling use case. Instead, we propose ExtraFerm as a unique tool that can be embedded in quantum-classical workflows to recover signal from noisy samples from large, application-scale quantum circuits.
@@
We apply ExtraFerm to the simulation of the local unitary cluster Jastrow (LUCJ) ansatz [motta2023lucj], which has recently been adopted for diverse applications in quantum simulation of chemical systems [blunt2025quantum, shajan2025toward, kaliakin2025implicit_solvent, liepuoniute2025quantumcentric, robledo2025sqd]. When mapped to a quantum circuit using the Jordan-Wigner transformation [ortiz2001quantum], the LUCJ ansatz decomposes into particle number-conserving matchgates and controlled-phase gates, making it amenable to simulation by ExtraFerm in certain regimes. To highlight the regimes in which ExtraFerm may be useful, we use LUCJ circuits to study the error and latency performance of ExtraFerm. Note that LUCJ circuits are flexible in the sense that repeated layers of them can be used to implement arbitrary particle number-conserving extended matchgate circuits, providing us with a generic but practically useful benchmark.
@@
Finally, as an end-to-end practical use case, we show how to use ExtraFerm to boost the performance of sample-based quantum diagonalization (SQD) [robledo2025sqd]. SQD is an extension of quantum-selected subspace interaction (QSCI) [kanno2023quantum], an algorithm that samples configurations from a quantum computer and uses them to select a subspace in which to diagonalize a molecular Hamiltonian. SQD adds an error mitigation procedure called configuration recovery that attempts to correct sampled bitstrings that were affected by noise and do not satisfy symmetries of the system. Configuration recovery greatly improves performance on noisy quantum processors, leading to the adoption of SQD for some of the largest demonstrations of quantum chemistry on quantum processors to date [liepuoniute2025quantumcentric, kaliakin2025implicit_solvent, shajan2025toward].
@@
    We demonstrate how to integrate our ExtraFerm with SQD, a hybrid quantum-classical energy estimation algorithm to improve the accuracy and variance of molecular ground-state energy calculations.
Figure 1: A high-level depiction of how ExtraFerm implements the Raw Estimate and Exact algorithms. The number of trajectories, $t$, can be given arbitrarily for Raw Estimate, while Exact sums over all possible trajectories. In practice, we parallelize this computation across bitstrings and then parallelize again across trajectories. Separate trajectories are generated for each bitstring to preserve the independence of Born-rule probability estimates.
@@ II-A Classical Simulation of Quantum Circuits
It is important to distinguish the very different computational tasks which are termed “simulation”. ExtraFerm computes additive-precision estimates of individual Born-rule probabilities and can be thought of as an approximate strong simulator in the sense of Ref. [jozsa2013classical]. Conversely, a weak simulator would sample from the output distribution of a quantum circuit. While it might seem intuitive that a strong simulator is stronger than a weak simulator, the precise comparison between the different simulation tasks is subtle and depends strongly on the error allowed [pashayan2020estimation]. However, our simulator can be used to approximately sample from the Born-rule distribution when combined with the Qubit-by-qubit algorithm [bravyi2022simulate] or the Gate-by-gate algorithm.
@@ II-B The Theory of Matchgates
Matchgates are a class of classically simulable quantum gates formalized by Valiant [valiant2002quantum]. Shortly afterwards, Terhal and DiVincenzo [terhal2002classical] discovered that matchgates are an equivalent expression of non-interacting fermions, giving these circuits a natural physical interpretation. The term ‘extended matchgate’ refers to circuits that are primarily composed of matchgates but include a limited quantity of non-matchgates—in our case, controlled-phase gates. Together, matchgates and controlled-phase gates allow for universal quantum computation [PhysRevA.73.042313, brod2011extending]. However, current matchgate-based chemistry ansatze are particle number-conserving, meaning that the Hamming weight of all measurement outcomes are the same as the initial Hamming weight of the system, as specified by Pauli-X gates. These are the circuits that ExtraFerm is designed for.
@@
where $a_{i}^{\dagger}$ denotes the $i^{\text{th}}$ creation operator defined via the Jordan-Wigner transformation. Note that while $M$ is a $2^{n}\times 2^{n}$ unitary matrix, $V$ is only $n\times n$. We refer to these ‘$V$’ matrices as mode transformation matrices. Let $\left\langle b\right|$ and $\left|a\right\rangle$ be two computational basis states with the same Hamming weights. Let $R$ be the indices of the bits of $\left\langle b\right|$ that are equal to 1 and let $C$ be the indices of the bits of $\left|a\right\rangle$ that are equal to 1.
@@
where $\widetilde{V}$ is the submatrix of $V_{\text{circuit}}$ obtained by selecting rows $r\in R$ and $c\in C$ [terhal2002classical].
@@ II-C Simulation of Extended Matchgate Circuits
ExtraFerm builds upon the recently proposed Algorithm 1 from Reardon-Smith et al. [reardon2024improved] and optimizes it for particle number-conserving extended matchgate circuits. We provide the details of this in section III-A. Simultaneously with [reardon2024improved], Diaz and Koenig [Dias2024classicalsimulation] developed alternative methods for extended matchgate simulation with essentially identical runtime scaling. Very recently, another approach has been developed with worse asymptotic scaling but significant improvements to the polynomial component of the runtime [wille2025classicalsimulationparitypreservingquantum]. These polynomial improvements may prove more relevant than the exponential component of the runtime in some parameter regimes.
@@ III-A Algorithms
ExtraFerm is capable of performing three different algorithms: Raw Estimate, Estimate, and Exact. Raw Estimate and Estimate, whose names follow the formalization in Ref. [pashayan2022fast], compute estimates of Born-rule probabilities, while Exact computes probabilities exactly. The Raw Estimate algorithm, which is depicted in Fig. 1, is the flagship capability of ExtraFerm. Given an additive error $\epsilon$, failure probability $\delta$, and probability upper bound $p_{\text{max}}$, Raw Estimate estimates the Born-rule probability $\hat{p}$ of a bitstring such that
@@
Raw Estimate does this by summing up contributions from trajectories, which are classically simulable approximations of the original circuit. Intuitively, each time we encounter a controlled-phase gate, we probabilistically branch between two possible paths. The number of trajectories that are needed to achieve a Born-rule probability estimate within $\epsilon$ for failure probability $\delta$ is exponential only in the extent $\xi^{\ast}$ of the circuit, given by the formula
@@ III-B Generating Trajectories
Figure 2: A visualization of how a trajectory $x$ and corresponding mode transformation matrix $V(x)$ is generated from a circuit. Each matchgate $m_{i}$ is unchanged while the controlled-phase gates are probabilistically assigned to either $d_{0}(\theta_{j})$ or $d_{0}(\theta_{j})$ based on $|\theta_{j}|$. The first controlled-phase gate is assigned to $d_{0}(\theta_{0})$ with probability 11 while the second controlled-phase gate is assigned to $d_{0}(\theta_{1})$ with probability $1/2$. Note that controlled-phase gates may operate on any two qubits and are not restricted to adjacent qubits.
@@
This calculations is the basis of the Exact algorithm (see Appendix section -A for derivation).
Allowing for a small additive error $\epsilon$ and failure probability $\delta$, Monte Carlo methods can be used to obtain an estimate for this Born-rule probability. Rather than summing up the amplitudes contributed by all $2^{k}$ trajectories, we can generate sufficiently many trajectories from the product probability distribution $P=\prod_{j=1}^{k}p(x_{j})$:
@@
To generate a trajectory, for every $\theta_{j}$, one can draw $u_{j}\sim\text{Unif}[0,1]$ and set
@@
  • computational basis states $\left\langle b\right|$ and $\left|a\right\rangle$
@@
  • computational basis states $\left\langle b\right|$ and $\left|a\right\rangle$
@@ IV-A Overview
We have open sourced ExtraFerm on Github, and it can be installed with pip install extraferm. ExtraFerm consists of a Python interface and Rust backend. The Python interface integrates directly with Qiskit [qiskit], while the Rust backend allows for multithreading of compiled binaries across multiple CPU cores, when available. Interaction between Python and Rust is facilitated by the PyO3 library, which allows Rust functions to be directly called directly from Python.
The machinery behind ExtraFerm’s Raw Estimate and Exact algorithms, depicted in Fig. 1, is doubly parallel. We use Rust’s Rayon library to parallelize the per-bitstring probability computations as well as each bitstring’s per-trajectory computations. A variety of other small optimizations led to further performance increases. For example, we represent states in a bit-packed format: Python uses native integers, while Rust leverages unsigned 128-bit integers (u128), enabling efficient bitwise operations. Additionally, we precompute and store values of $\sin(\theta/4)$ and $\cos(\theta/4)$ for all $\theta$, which reduces the number of computations inside hot loops.

### v2

@@ Abstract
We present and open source a quantum circuit simulator tailored to chemistry applications. More specifically, our simulator can compute the Born-rule probabilities of samples obtained from circuits containing passive fermionic linear optical elements and controlled-phase gates. We support both approximate and exact calculation of probabilities, and for approximate probability calculation, our simulator’s runtime is exponential only in the magnitudes of the circuit’s controlled-phase gate angles. This makes our simulator useful for simulating certain systems that are beyond the reach of conventional state vector methods. We demonstrate our simulator’s utility by simulating the local cluster unitary Jastrow (LUCJ) ansatz and integrating it with sample-based quantum diagonalization (SQD) to improve the accuracy of molecular ground-state energy estimates. Applied to a 52-qubit N2 system, we observe accuracy improvements of up to 46% over the baseline SQD implementation with negligible computational overhead. More generally, we highlight a regime in which our simulator achieves substantially superior latency scaling and exponentially superior memory scaling over a tensor network simulator and a state vector simulator. As an efficient and flexible tool for simulating quantum chemistry circuits, our simulator enables new opportunities for enhancing near-term quantum algorithms in chemistry and related domains.
@@ I Introduction
Among the most promising near-term applications of quantum computing is quantum computational chemistry [cao2019quantum, mcardle2020quantum], which has attracted considerable attention due to potentially transformative applications in drug development [santagati2024drug] and materials science [alexeev2024quantum], among other fields. Despite the excitement around this application, hardware demonstrations have been limited by the noise present in current quantum processors. Quantum error correction [gottesman1997stabilizer, fowler2012surface, shor1995scheme, steane1996error, calderbank1996good, kitaev2003fault] offers a long-term solution to faulty hardware, but a fully fault-tolerant quantum computer has yet to be built. In the near-term, quantum error mitigation techniques [temme2017error, li2017efficient, huggins2021virtual, liao2024machine, cincio2021machine] can improve results obtained from noisy quantum processors, in which many of these techniques involve classical simulation of quantum circuits [czarnik2021error, filippov2023scalabletensornetworkerrormitigation, eddins2024lightcone]. Because classical simulation of quantum circuits is intractable in general, these simulations must introduce approximations, or are exact only when applied to specific narrow classes of circuits. Simulation methods tailored to circuits relevant to quantum chemistry may enable new techniques for enhancing the performance of chemistry workloads on quantum computers.
Figure 1: The green star indicates the class of circuits that ExtraFerm is designed to address: circuits containing $m$ matchgates and $k$ controlled-phase gates where $k\ll m$. This work concentrates on particle number-conserving circuits. Note that matchgates + controlled-phase gates form a universal gate set.
@@
In this work, we introduce ExtraFerm [^1] https://github.com/zhassman/ExtraFerm, an open-source quantum circuit simulator that can be used to compute the Born-rule probabilities of samples (bitstrings) from circuits composed of passive fermionic linear optical elements, also known as particle number-conserving matchgates, and controlled-phase gates. These circuits are highly relevant to the field of quantum computational chemistry for simulating quantum many-body systems. Fig. 1 shows where these circuits sit relative to other well-studied gate families. ExtraFerm can perform both exact and approximate probability calculations for measurements in the output distributions of these circuits. For approximate calculation, the only exponential component of ExtraFerm’s runtime is a linear dependence in the extent of the quantum circuit it is simulating. The extent is larger for controlled phase gates with larger phases and is multiplicative as more controlled phase gates are added (see Refs. [reardon2024improved, Dias2024classicalsimulation, reardonsmith2024fermioniclinearopticalextent]). For exact probability calculation, ExtraFerm’s runtime is exponential only in the number of controlled-phase gates.
@@
Notably, this means that for both approximate and exact calculation, ExtraFerm’s performance is polynomial in the number of qubits and matchgates. Unlike conventional state vector simulation, which tracks the amplitudes of all bitstrings, ExtraFerm computes Born-rule probabilities for a pre-specified subset of the output distribution. Indeed, for small systems, one can use ExtraFerm to simulate the entire support of a circuit, but we do not expect this to be ExtraFerm’s most compelling use case. Instead, we propose ExtraFerm as a unique tool that can be embedded in quantum-classical workflows to recover signal from noisy samples from large, application-scale quantum circuits.
@@
We apply ExtraFerm to the simulation of the local unitary cluster Jastrow (LUCJ) ansatz [motta2023lucj], which has recently been adopted for diverse applications in quantum simulation of chemical systems [blunt2025quantum, shajan2025toward, kaliakin2025implicit_solvent, liepuoniute2025quantumcentric, robledo2025sqd]. When mapped to a quantum circuit using the Jordan-Wigner transformation [ortiz2001quantum], the LUCJ ansatz decomposes into particle number-conserving matchgates and controlled-phase gates, making it amenable to simulation by ExtraFerm in certain regimes. To highlight the regimes in which ExtraFerm may be useful, we use LUCJ circuits to study the error and latency performance of ExtraFerm. Note that LUCJ circuits are flexible in the sense that repeated layers of them can be used to implement arbitrary particle number-conserving extended matchgate circuits, providing us with a generic but practically useful benchmark. Beyond this, Section V-A show that ExtraFerm offers significantly lower latency and exponentially better memory performance compared to both a tensor network and state vector simulator.
@@
Finally, as an end-to-end practical use case, we show how to use ExtraFerm to boost the performance of sample-based quantum diagonalization (SQD) [robledo2025sqd]. SQD is an extension of quantum-selected configuration interaction (QSCI) [kanno2023quantum], an algorithm that samples configurations from a quantum computer and uses them to select a subspace in which to diagonalize a molecular Hamiltonian. SQD adds an error mitigation procedure called configuration recovery that attempts to correct sampled bitstrings that were affected by noise and do not satisfy symmetries of the system. Configuration recovery greatly improves performance on noisy quantum processors, leading to the adoption of SQD for some of the largest demonstrations of quantum chemistry on quantum processors to date [liepuoniute2025quantumcentric, kaliakin2025implicit_solvent, shajan2025toward].
@@
    We demonstrate how to integrate ExtraFerm with SQD, a hybrid quantum-classical energy estimation algorithm to improve the accuracy and variance of molecular ground-state energy calculations.
Figure 2: A high-level depiction of how ExtraFerm computes the Born-rule probability of a bitstring using Raw Estimate or Exact. The number of trajectories, $t$, can be given arbitrarily for Raw Estimate, while Exact sums over all possible trajectories. In practice, we parallelize this computation across bitstrings and then parallelize again across trajectories. Separate trajectories are generated for each bitstring to preserve the independence of Born-rule probability estimates.
@@ II-A Classical Simulation of Quantum Circuits
It is important to distinguish the very different computational tasks which are termed “simulation”. ExtraFerm computes additive-precision estimates of individual Born-rule probabilities and can be thought of as an approximate strong simulator in the sense of Ref. [jozsa2013classical]. Conversely, a weak simulator would sample from the output distribution of a quantum circuit. While it might seem intuitive that a strong simulator is stronger than a weak simulator, the precise comparison between the different simulation tasks is subtle and depends on the error allowed [pashayan2020estimation]. However, our simulator can be used to approximately sample from the output distribution of a circuit when combined with the Qubit-by-qubit algorithm [bravyi2022simulate] or the Gate-by-gate algorithm.
@@ II-B The Theory of Matchgates
Matchgates are a class of classically simulable quantum gates formalized by Valiant [valiant2002quantum]. Shortly afterwards, Terhal and DiVincenzo [terhal2002classical] discovered that matchgates are an equivalent expression of non-interacting fermions, giving these circuits a natural physical interpretation. The term ‘extended matchgate’ refers to circuits that are primarily composed of matchgates but include a limited quantity of non-matchgates—in our case, controlled-phase gates. Together, matchgates and controlled-phase gates allow for universal quantum computation [PhysRevA.73.042313, brod2011extending]. However, most matchgate-based chemistry ansatze are particle number-conserving, meaning that the Hamming weights of all measurement outcomes are the same as the initial Hamming weight of the system, as specified by Pauli-X gates. These are the circuits that ExtraFerm is designed for.
@@
where $a_{i}^{\dagger}$ denotes the $i^{\text{th}}$ creation operator defined via the Jordan-Wigner transformation. Note that while $M$ is a $2^{n}\times 2^{n}$ unitary matrix, $V$ is only $n\times n$. We refer to these ‘$V$’ matrices as mode transformation matrices. Let $\left|a\right\rangle$ and $\left|b\right\rangle$ be two computational basis states with the same Hamming weights. Let $R$ be the indices of the bits of $\left|b\right\rangle$ that are equal to 1 and let $C$ be the indices of the bits of $\left|a\right\rangle$ that are equal to 1. Then
@@
where $\widetilde{V}$ is the submatrix of $V$ obtained by selecting rows $r\in R$ and $c\in C$ [terhal2002classical].
@@ II-C Simulation of Extended Matchgate Circuits
ExtraFerm builds upon the recently proposed mathematical framework from Reardon-Smith [reardon2024improved] and optimizes it for particle number-conserving circuits. We provide the details of this in section III-A. Simultaneously with [reardon2024improved], Diaz and Koenig [Dias2024classicalsimulation] developed alternative methods for extended matchgate simulation with essentially identical runtime scaling. Very recently, another approach has been developed with worse asymptotic scaling but significant improvements to the polynomial component of the runtime [wille2025classicalsimulationparitypreservingquantum]. These polynomial improvements may prove more relevant than the exponential component of the runtime in some parameter regimes.
@@ III-A Algorithms
ExtraFerm is capable of performing three different algorithms: Raw Estimate (Algorithm 1), Estimate (Algorithm 2), and Exact. Raw Estimate and Estimate, whose names follow the formalization in Ref. [pashayan2022fast], compute estimates of Born-rule probabilities, while Exact computes probabilities exactly. The Raw Estimate algorithm, which is depicted in Fig. 2, is the flagship capability of ExtraFerm. Given an additive error $\epsilon$, failure probability $\delta$, and probability upper bound $p_{\text{max}}$, Raw Estimate estimates the Born-rule probability $\hat{p}$ of a bitstring such that
@@
Raw Estimate does this by summing up contributions from trajectories, which are classically simulable matchgate approximations of the original circuit. Intuitively, each time we encounter a controlled-phase gate, we probabilistically branch between two possible paths. The number of trajectories that are needed to achieve a Born-rule probability estimate within $\epsilon$ for failure probability $\delta$ is exponential only in the extent $\xi^{\ast}$ of the circuit, given by the formula
@@
Summary: ExtraFerm has the ability to perform three simulation algorithms, each with different use cases.
@@ III-B Generating Trajectories
Figure 3: A visualization of how a trajectory $x$ and corresponding mode transformation matrix $V(x)$ are generated from a circuit. Each matchgate $m_{i}$ is unchanged while the controlled-phase gates are probabilistically assigned to either $d_{0}(\theta_{j})$ or $d_{1}(\theta_{j})$ based on $|\theta_{j}|$. In the above example, the trajectory $x=010$ has been sampled. The first controlled-phase gate is assigned to $d_{0}(\theta_{0})$ with probability 11 while the second and third controlled-phase gates are assigned to $d_{1}(\theta_{1})$ and $d_{0}(\theta_{2})$, each with probability 1/2. Note that controlled-phase gates may operate on any two qubits; they are not restricted to nearest neighbors.
@@
This calculation forms the basis of the Exact algorithm.
Allowing for a small additive error $\epsilon$ and failure probability $\delta$, Monte Carlo methods can be used to obtain an estimate for this Born-rule probability. Rather than summing up the amplitudes contributed by all $2^{k}$ trajectories, we can sample trajectories from the product probability distribution $P=\prod_{j=1}^{k}p(x_{j})$:
@@
To generate a trajectory, for every controlled-phase gate angle $\theta_{j}$, one can draw $u_{j}\sim\text{Unif}[0,1]$ and set
@@
Summary: For exact simulation, we sum over all possible matchgate approximations of the circuit, whereas for the approximate case we sum over a probabilistically generated multiset of them.
@@
  • computational basis states $\left|a\right\rangle$ and $\left|b\right\rangle$
@@
  • computational basis states $\left|a\right\rangle$ and $\left|b\right\rangle$
@@ IV-A Overview
We have open sourced ExtraFerm on Github, and it can be installed with pip install extraferm. ExtraFerm consists of a Python interface and Rust backend. The Python interface integrates directly with Qiskit [qiskit], while the Rust backend allows for multithreading of compiled binaries across multiple CPU cores, when available. Interaction between Python and Rust is facilitated by the PyO3 library, which allows Rust functions to be called directly from Python.
The machinery behind ExtraFerm’s Raw Estimate and Exact algorithms, depicted in Fig. 2, is doubly parallel. We use Rust’s Rayon library to parallelize the per-bitstring probability computations as well as each bitstring’s per-trajectory computations. A variety of other small optimizations led to further performance increases. For example, we represent states in a bit-packed format: Python uses native integers, while Rust leverages unsigned 128-bit integers (u128), enabling efficient bitwise operations. Additionally, we precompute and store values of $\sin(\theta_{j}/4)$ and $\cos(\theta_{j}/4)$ for all $\theta_{j}$, which reduces the number of computations inside hot loops.